OpenStudy (giitter):

Which operation should be performed first according to the order of operations? 87 – 25 + 39 • 8 + 31 + 10 A. 31 + 10 B. 87 – 25 C. 25 + 39 D. 39 • 8

Awolflover1 (awolflover1):

P- ( ) E-2^3 M- x D-÷ A- + S- -

Awolflover1 (awolflover1):

okay so they don't have parenthesis? right?

Awolflover1 (awolflover1):

87 – 25 + 39 • 8 + 31 + 10 Lets go down the list

OpenStudy (giitter):

ok

Awolflover1 (awolflover1):

so no exponents right?

OpenStudy (giitter):

right

Awolflover1 (awolflover1):

alright what about times? we have that right?

OpenStudy (giitter):

Right,

Awolflover1 (awolflover1):

yes there we go!

OpenStudy (giitter):

to we would multiply first?

Awolflover1 (awolflover1):

Yep!

OpenStudy (giitter):

Okeh

Awolflover1 (awolflover1):

Cause we don't see any exponents or parenthesis so yep!

OpenStudy (giitter):

Is the answer to that \[312?\]

OpenStudy (giitter):

and then we would add the other part? That's the part that confuses me.

OpenStudy (giitter):

i tried my calculator twice

Awolflover1 (awolflover1):

Okay >.< did you try \[87-25+25+39\times8+31+10\] 87-25=62 and 39 x 8=? and 31+10=?

Awolflover1 (awolflover1):

\[62+312+41=\]

OpenStudy (giitter):

I got it :3

Awolflover1 (awolflover1):

yay!